# Addressing the Rank Degeneration in Sequential Recommendation via Singular Spectrum Smoothing

### v1

@@ 4.5. Optimization with Dual Smoothing
With singular spectrum smoothing on both sequence output embeddings side $\mathcal{L}_{\text{seq}}$ and item side $\mathcal{L}_{\text{ite}}$, the final optimization loss $\mathcal{L}$ with the recommendation loss in Eq. (5) to be minimized is defined as follows:
@@ 5. Experiments
  • RQ2: Does SPMRec outperform existing degeneration methods?
  • RQ3: How does each component of SPMRec contribute to the performance?
  • RQ4: Does SPMRec alleviate the rank collapse problem of Transformer?
  • RQ5: How does our item side spectrum smoothing $\beta$ affect the diversity and recommendation performance?
  • RQ6: Where do improvements of SPMRec come from?
5.1. Datasets
We present the statistics of the dataset in Table 1. We evaluate all models in four benchmarks public Amazon review datasets [^1] http://deepyeti.ucsd.edu/jianmo/amazon/index.html. In the Amazon reviews dataset, several categories of user-item interactions with timestamps are available. We choose Beauty, Toys and Games (Toys), Tools and Home (Tools), and Office Products (Office) categories in our experiments as these four categories are widely adopted benchmark categories (Tang and Wang, 2018; Kang and McAuley, 2018; Sun et al., 2019; Fan et al., 2021; Li et al., 2020; Fan et al., 2022). We treat the existence of user-item reviews as positive user-item interactions. For each user, we sort the interacted items chronologically to form the user interaction sequence. The last interaction of the user sequence is used for testing and the second to last one for validation. We adopt the standard 5-core pre-processing setting on users (Tang and Wang, 2018; Kang and McAuley, 2018; Sun et al., 2019; Fan et al., 2021; Li et al., 2020; Fan et al., 2022) to filter out users with less than five interactions. The detailed dataset statistics are presented in Table 1.
@@ 5.1. Datasets
Total 63K 46K 554K - -
@@ 5.2. Baselines and Hyper-parameter Grid Search
We implement SPMRec with Pytorch based on the S3-Rec (Zhou et al., 2020) code base. We grid search all hyper-parameters and report the test performance based on the best validation MRR over all combinations of hyper-parameters experiments. We search the representation dimension in $\{64,128\}$, maximum sequence length from $\{50,100\}$, learning rate in $\{10^{-3},10^{-4}\}$, the $L2$ regularization control weight from $\{0.0,10^{-1},10^{-2},10^{-3}\}$, dropout rate from $\{0.3,0.5,0.7\}$ for all methods. For Transformer-based methods, we search the number of layers from $\{1,2,3\}$, and the number of heads in $\{1,2,4\}$. We adopt the early stopping strategy when the validation MRR does not increase for 50 epochs.
5.3. Evaluation
We predict the next item and generate the top-N recommendation list for each user based on the dot-product between sequence output embedding and the entire item set in descending order. We rank all items for all models so that no item sampling bias is introduced in evaluation for fair comparisons (Krichene and Rendle, 2020). The evaluation includes classical top-N ranking metrics, Recall@N and NDCG@N. Recall@N measures the percentage of sequences being recommended true preferred item in the top-N list. NDCG@N weights the ranking position of the ground truth preferred item in the ranking list by giving exponentially smaller weights to lower-ranked items. We present the averaged results over all test users, which is also the entire user set. The test results are reported based on the best validation results. We report metrics in multiple Ns, including $N=\{5,10,40\}$, which are widely adopted in most existing SR methods (Kang and McAuley, 2018; Sun et al., 2019; Fan et al., 2022).
We also evaluate the diversity of the recommendation list for all test users, which is measured as the Intra-list diversity (Zhang and Hurley, 2008; Ziegler et al., 2005) for each user’s top-N recommendation list. We focus on the top-10 ranking list for diversity calculation. The intra-list diversity is calculated as follows:
(14) $Diversity(u)=1-\frac{\sum_{i\in\mathcal{I}_{u}^{N}}\sum_{j\in\mathcal{I}_{u}^{N}}\text{cos}(\mathbf{E}_{i},\mathbf{E}_{j})}{|\mathcal{I}_{u}^{N}|*|\mathcal{I}_{u}^{N}|},$
where $\mathcal{I}_{u}^{N}$ denotes the top-$N$ recommendation list for the user $u$, $\text{cos}(\cdot,\cdot)$ represents the cosine similarity, $\mathbf{E}$ is the item embedding table defined in Section 3.2, and the reported diversity is the average diversity over all users. The larger value of diversity indicates higher intra-list diversity for the recommendation.
@@ 5.4. Overall Comparisons (RQ1)
We present the overall performance comparison with relevant state-of-the-art baselines in Table 2. From Table 2, we can easily conclude the superiority of SPMRec in providing sequential recommendations in various datasets and metrics. We have several observations to conclude as follows:
@@
  • SPMRec achieves the best performance and outperforms SASRec with large improvement margins. In most metrics and all datasets, SPMRec significantly outperforms the second-best baseline with 0.52% to 12.33%. The improvements demonstrate the effectiveness of SPMRec in the sequential recommendation. Furthermore, as SPMRec is built on top of the original implementation of SASRec (Kang and McAuley, 2018), we observe 6.92% to 22.53% relative improvements over SASRec with both sequence output and item embeddings spectrum smoothings. We argue that the improvements originate from the singular spectrum smoothings on both sequence and item sides.
  • SPMRec can better alleviate the sequence rank collapse issue. Comparing the AUSC of SPMRec and baselines, we can observe that SPMRec consistently achieves better AUSC values, indicating the better capability of alleviating the sequence rank issue in SR.
@@
  • SPMRec achieves larger improvements in sparser datasets. Comparing SPMRec and SASRec, we can observe at least 17% improvements in both Beauty and Tools categories datasets, which have lower densities of 0.05% and 0.07%, respectively. However, in Toys and Office datasets with higher densities of 0.08% and 0.44%, the improvements are at most 14%. This observation demonstrates that the proposed singular spectrum smoothing SPMRec benefits sparser datasets, especially for datasets with a large portion of short sequences.
@@
  • The consideration of Frobenius norm benefits the recommendation. STOSA adopts the sum of two L2 distances on mean and covariance embeddings and DuoRec utilizes the uniformity loss based on the exponential L2 distance. We can observe that STOSA and DuoRec both outperform the dot-product-based SASRec framework. Unlike the dot product approach, L2 distance considers the Frobenius norm of embeddings, even if it does not constrain the Frobenius norm, which again is the upper bound of the largest singular value of embeddings. This observation further demonstrates the need to consider the Frobenius norm in embedding learning. Moreover, SPMRec performs better than STOSA and DuoRec, demonstrating the necessity of considering the regularization from the singular spectrum perspective.
@@
5.5. Degeneration Methods Comparison (RQ2)
@@
(15) $\mathcal{L}_{\text{item}}=\frac{1}{|\mathcal{V}|\times|\mathcal{V}|}\sum_{i=1}^{|\mathcal{V}|}\sum_{j=1}^{|\mathcal{V}|}\frac{\mathbf{E}_{i}^{\top}\mathbf{E}_{j}}{||\mathbf{E}_{i}||*||\mathbf{E}_{j}||},$
@@
(16) $\mathcal{L}_{\text{item}}=-\frac{1}{|\mathcal{V}|\times|\mathcal{V}|}\sum_{i=1}^{|\mathcal{V}|}\sum_{j=1}^{|\mathcal{V}|}||\mathbf{E}_{i}-\mathbf{E}_{j}||_{2}=-\frac{1}{|\mathcal{V}|\times|\mathcal{V}|}\sum_{i=1}^{|\mathcal{V}|}\sum_{j=1}^{|\mathcal{V}|}||\mathbf{E}_{i}||_{2}-2\mathbf{E}_{i}^{\top}\mathbf{E}_{j}+||\mathbf{E}_{j}||_{2}.$
Following our motivations in both sequence and item sides, we apply the cosine and Euclidean regularization methods to both sides. We report the best grid searched results of all regularization methods in Table 3. Note that the uniformity property adopted by DuoRec and DirectAU (Qiu et al., 2022; Wang et al., 2022) falls into the Euclidean methods category.
As shown in Table 3, neither cosine regularization or Euclidean regularization can consistently outperform the original SASRec without any regularization. Arguably, the potential reason is that neither of these two methods constrains the Frobenius norm of embeddings, which is the upper bound of the largest singular value, as discussed in Section 4.1. When cosine regularization and the Euclidean method are minimized, the naive solution is $||\mathbf{E}_{i}||_{2}\rightarrow\infty$ for all $i$ if we do not consider the supervised recommendation loss. Moreover, different from the word embedding task, sequences in the recommendation task are shorter on average, posing a greater challenge in addressing the degeneration and the rank collapse problem in SR. However, our proposed singular spectrum smoothing method SPMRec minimizes the Frobenius norm while maximizing the nuclear norm, achieving the goals of smoothing the singular value curve and suppressing the largest singular value simultaneously.
@@ 5.6.2. Item Embedding Smoothing Weight $\beta$
(a) Beauty
(b) Toys
(c) Tools
(d) Office
5.7. Sequential User Singular Values Decay Analysis (RQ4)
We visualize the singular values curve of sequence output embedding for our proposed SPMRec and SASRec, as shown in Fig. (5). We can observe that the proposed SPMRec has a smoother decaying curve than SASRec. Moreover, for SPMRec, the degrees of spectrum smoothness for all four datasets are 26.34 for Beauty, 20.76 for Tools, 20.25 for Toys, and 19.24 for Office. However, SASRec has lower degrees for spectrum smoothness as 19.54 for Beauty, 17.10 for Tools, 17.00 for Toys, and 13.67 for Office. SPMRec consistently has larger degrees of spectrum smoothness than SASRec, which demonstrates the effectiveness of SPMRec in preventing Transformer from rank collapse for the sequential recommendation.
(a) Beauty
(b) Toys
(c) Tools
(d) Office
Figure 6. Relationship between item spectrum smoothing weight $\beta$ and recommendation diversity. Best Viewed in colors.
5.8. Diversity and Item Spectrum Smoothing (RQ5)
Item Spectrum Smoothing controlled by $\beta$ is closely related to the recommendation diversity. Moreover, the recommendation performance drops when we demand a high recommendation diversity. We demonstrate a strong connection between our item spectrum smoothing and the recommendation diversity by showing the relationship of $\beta$ and the diversity, as shown in Fig. (6). In Fig. (6), the x-axis denotes the value of $\beta$. As the $\beta$ becomes larger, we can observe that the diversity consistently increases, in all four datasets, demonstrating the strong connection between the recommendation diversity and our proposed item spectrum smoothing component. Another observation is that the NDCG@10 drops significantly when we enforce too large $\beta$, which also indicates a higher diversity. This observation demonstrates that a proper selection of $\beta$ can achieve satisfactory recommendation performance and diversity simultaneously. However, demanding higher diversity can hurt the recommendation performance.
@@ 5.8. Diversity and Item Spectrum Smoothing (RQ5)
5.9. Improvements Analysis (RQ6)
In this section, we investigate the improvements of SPMRec from both sequence and item sides for a better understanding of the advantages of SPMRec. The improvements analysis are shown in Fig. (7) and Fig. (8), for sequences and items respectively. We compare with three strong baselines, including SASRec, STOSA, and DuoRec (having data augmentations).
5.9.1. Improvements w.r.t. Sequence Length
SPMRec is highly effective in short sequences recommendation. We group sequences (users) based on the number of training interactions (sequence length) of users. Within each group, we average the testing performances of sequences in the group. The comparisons on each group are shown in Fig. (7). We can observe that SPMRec achieves significant improvements in short sequences. Specifically, the improvements over the best STOSA range from 7.21% to 23.04%. The recommendation for short sequences is challenging due to limited interaction signals, which is known as the cold start problem. SPMRec regularizes the sequence output embeddings and prevents the rank collapse so that short sequences can be distinguishable and are properly modeled. We can also observe that SPMRec has inferior performances in the longest sequence group, which is even worse than SASRec in the Toys dataset. However, the short sequences are in majority in both the used datasets and practical applications. With the proposed singular spectrum smoothing, SPMRec benefits short sequences with the sacrifices of long sequences.
5.9.2. Improvements w.r.t. Item Popularity
The SPMRec focuses the representation learning for popular items with the best searched $\beta$. We group items based on the number of training reviews of items, i.e., item popularity. We report the average testing performance of items in each group. We show the comparisons in Fig. (8). We can observe that SPMRec achieves significant improvements in the most popular groups of items while only comparative and even inferior performances in less popular groups. Fig. (8) shows the group performance with the best searched $\beta$, which emphasizes popular items. However, this observation does not contradict the motivation of item embedding singular spectrum smoothing as we observe that a large $\beta$ provides more balanced performances in different item groups while inferior overall sequences recommendation performances.

### v2

@@ 4.3. Optimization with Dual Smoothing
With singular spectrum smoothing on both sequence output embeddings side $\mathcal{L}_{\text{seq}}$ and item side $\mathcal{L}_{\text{item}}$, the final optimization loss $\mathcal{L}$ with the recommendation loss in Eq. (3) to be minimized is defined as follows:
@@ 5. Experiments
  • RQ2: How does our item side spectrum smoothing $\beta$ affect the diversity and recommendation performance?
  • RQ3: Does SPMRec outperform existing degeneration solutions?
  • RQ4: How does each component of SPMRec contribute to the performance?
  • RQ5: Where do improvements of SPMRec come from?
@@
Total 129K 74K 1.10M - -
5.1. Datasets
We present the statistics of the dataset in Table 1. We evaluate all models in four benchmarks public Amazon review datasets [^1] http://deepyeti.ucsd.edu/jianmo/amazon/index.html. We choose Beauty, Toys and Games (Toys), Tools and Home (Tools), and Office Products (Office) categories in our experiments as these four categories are widely adopted benchmark categories (Tang and Wang, 2018; Kang and McAuley, 2018; Sun et al., 2019; Fan et al., 2021; Li et al., 2020; Fan et al., 2022). We treat the existence of user-item reviews as positive user-item interactions. For each user, we sort the interacted items chronologically to form the user interaction sequence. The last interaction of the user sequence is used for testing and the second to last one for validation. We adopt the standard 5-core pre-processing setting on users (Tang and Wang, 2018; Kang and McAuley, 2018; Sun et al., 2019; Fan et al., 2021; Li et al., 2020; Fan et al., 2022) to filter out users with less than five interactions. The detailed dataset statistics are presented in Table 1.
@@ 5.2. Baselines and Hyper-parameter Grid Search
5.3. All Items Ranking and Diversity Evaluation
We generate the top-N recommendation list for each user based on the dot-product between sequence output embedding and the entire item set in descending order. We rank all items for all models so that no item sampling bias is introduced in evaluation for fair comparisons (Krichene and Rendle, 2020). The evaluation includes classical top-N ranking metrics, Recall@N and NDCG@N. We present the averaged results over all test users. The test results are reported based on the best validation results. We report metrics in multiple Ns, including $N=\{5,10,40\}$, which are widely adopted in most existing SR methods (Kang and McAuley, 2018; Sun et al., 2019; Fan et al., 2022).
We evaluate the diversity of the recommendation list for all test users in two widely adopted metrics, including the Intra-list diversity (Zhang and Hurley, 2008; Ziegler et al., 2005) and the Coverage@100 of item categories for each user’s top-N recommendation list. We focus on the top-10 ranking list for intra-list diversity calculation and top-100 for the coverage metric. The intra-list diversity is calculated as follows:
(15) $Diversity(u)=1-\frac{\sum_{i\in\mathcal{I}_{u}^{N}}\sum_{j\in\mathcal{I}_{u}^{N}}\text{cos}(\mathbf{M}_{i},\mathbf{M}_{j})}{|\mathcal{I}_{u}^{N}|*|\mathcal{I}_{u}^{N}|},$
where $\mathcal{I}_{u}^{N}$ denotes the top-$N$ recommendation list for the user $u$, $\text{cos}(\cdot,\cdot)$ represents the cosine similarity, $\mathbf{M}$ is the item embedding table defined in Section 2.2, and the reported diversity is the average diversity over all users. The larger value of diversity indicates higher intra-list diversity for the recommendation. Note that, we also found similar improvements in other larger $N$s, such as $N=\{20,40\}$.
@@ 5.4. Overall Comparisons (RQ1)
We present the overall performance comparison with relevant state-of-the-art baselines in Table 2. From Table 2, we can easily conclude the superiority of SPMRec in providing sequential recommendations in various datasets and achieving not only better performance but also more diversity. We have several observations to conclude as follows:
@@
  • SPMRec achieves the best performance and diversity, and outperforms SASRec with large improvement margins. In most performance metrics and all datasets, SPMRec significantly outperforms the second-best baseline with 0.52% to 12.33% in ranking performance and 4.71% to 14.41% in diversity. The improvements demonstrate the effectiveness of SPMRec in the sequential recommendation, while the recommendation diversity is also noticeably improved. Furthermore, as SPMRec is built on top of the original implementation of SASRec (Kang and McAuley, 2018), we observe 6.92% to 22.53% relative improvements over SASRec. We argue that the improvements originate from the singular spectrum smoothings on both sequence and item sides.
  • SPMRec can achieve better balance between recommendation quality and diversity. Comparing the diversity of SPMRec and baselines, we can observe that SPMRec consistently achieves better diversity results while having better performances, indicating the better capability of SPMRec in balancing the recommendation quality and diversity.
@@
  • SPMRec achieves performance larger improvements in sparser datasets. Comparing SPMRec and SASRec, we can observe at least 17% improvements in both Beauty and Tools categories datasets, which have lower densities of 0.05% and 0.07%, respectively. The recommendation diversity is also significantly improved, especially for tools with a 14.43% improvement. However, in Toys and Office datasets with higher densities of 0.08% and 0.44%, the improvements are at most 14%. This observation demonstrates that the proposed singular spectrum smoothing SPMRec benefits sparser datasets, especially for datasets with a large portion of short sequences.
@@
  • The consideration of Frobenius norm benefits the recommendation and diversity. STOSA adopts the sum of two L2 distances on mean and covariance embeddings and DuoRec utilizes the uniformity loss based on the exponential L2 distance. We can observe that STOSA and DuoRec both outperform the dot-product-based SASRec framework in recommendation quality and diversity. This observation further demonstrates the need to consider the Frobenius norm in embedding learning. Moreover, SPMRec performs better than STOSA and DuoRec, demonstrating the necessity of considering the regularization from the singular spectrum perspective.
(a) Beauty
(b) Toys
(c) Tools
(d) Office
Figure 4. Relationship between item spectrum smoothing weight $\beta$ and diversity based on embeddings.
(a) Beauty
(b) Toys
(c) Tools
(d) Office
Figure 5. Relationship between item spectrum smoothing weight $\beta$ and diversity based on coverage@100.
5.5. Diversity and Item Smoothing $\mathcal{L}_{\text{item}}$ (RQ2)
Item Spectrum Smoothing controlled by $\beta$ is closely related to the recommendation diversity. We demonstrate a strong connection between our item spectrum smoothing and the recommendation diversity by showing the relationship of $\beta$ and the diversity, as shown in Fig. (4). In Fig. (4), the x-axis denotes the value of $\beta$. As the $\beta$ becomes larger, we can observe that the diversity consistently increases, in all four datasets, demonstrating the strong connection between the recommendation diversity and our proposed item spectrum smoothing component.
Moreover, the recommendation performance drops when we demand too high recommendation diversity. Another observation is that the NDCG@10 drops significantly when we set $\beta$ to a too-large value, which also indicates a higher diversity. This observation demonstrates that a proper selection of $\beta$ can achieve satisfactory recommendation performance and diversity simultaneously. However, demanding higher diversity can hurt the recommendation performance.
@@ 5.5. Diversity and Item Smoothing $\mathcal{L}_{\text{item}}$ (RQ2)
5.6. Degeneration Methods Comparison (RQ3)
@@
(16) $\mathcal{L}_{\text{item}}=\frac{1}{|\mathcal{V}|\times|\mathcal{V}|}\sum_{i=1}^{|\mathcal{V}|}\sum_{j=1}^{|\mathcal{V}|}\frac{\mathbf{M}_{i}^{\top}\mathbf{M}_{j}}{||\mathbf{M}_{i}||\times||\mathbf{M}_{j}||},$
@@
(17) $\mathcal{L}_{\text{item}}$ $=-\frac{1}{|\mathcal{V}|\times|\mathcal{V}|}\sum_{i=1}^{|\mathcal{V}|}\sum_{j=1}^{|\mathcal{V}|}||\mathbf{M}_{i}-\mathbf{M}_{j}||_{2}.$
We apply the cosine and Euclidean regularization methods to both sides. We report the best grid searched results of all regularization methods in Table 3. Note that the uniformity property adopted by DuoRec and DirectAU (Qiu et al., 2022; Wang et al., 2022b) falls into the Euclidean methods category.
As shown in Table 3, neither cosine regularization or Euclidean regularization can consistently outperform the original SASRec without any regularization. Arguably, the potential reason is that neither of these two methods constrains the Frobenius norm of embeddings, which is the upper bound of the largest singular value, as discussed in Section 3.1. When cosine regularization and the Euclidean method are minimized, the naive solution is $||\mathbf{E}_{i}||_{2}\rightarrow\infty$ for all $i$ if we do not consider the supervised recommendation loss. Moreover, different from the word embedding task, sequences in the recommendation task are shorter on average, posing a greater challenge in addressing the degeneration in SR. However, our proposed singular spectrum smoothing method SPMRec minimizes the Frobenius norm while maximizing the nuclear norm, achieving the goals of smoothing the singular value curve and suppressing the largest singular value simultaneously.
@@ 5.7.2. Item Embedding Smoothing Weight $\beta$
5.8. Improvements Analysis (RQ5)
5.8.1. Improvements w.r.t. Sequence Length
SPMRec is highly effective in short sequences recommendation. We group sequences (users) based on the number of training interactions (sequence length) of users. Within each group, we average the testing performances of sequences in the group. The comparisons on each group are shown in Fig. (8). We can observe that SPMRec achieves significant improvements in short sequences. Specifically, the improvements over the best STOSA range from 7.21% to 23.04%. The recommendation for short sequences is challenging due to limited interaction signals, which is known as the cold start problem. SPMRec regularizes the sequence output embeddings and prevents the degeneration so that short sequences can be distinguishable and are properly modeled. We can also observe that SPMRec has inferior performances in the longest sequence group, which is even worse than SASRec in the Toys dataset. However, the short sequences are in majority in both the used datasets and practical applications. With the proposed singular spectrum smoothing, SPMRec benefits short sequences with the sacrifices of long sequences.
@@ 5.8. Improvements Analysis (RQ5)
(a) Beauty
(b) Toys
(c) Tools
(d) Office
5.9. Sequential User Degeneration Analysis (RQ5)
We visualize the singular values curve of sequence output embedding for our proposed SPMRec and SASRec, as shown in Fig. (10). We can observe that the proposed SPMRec has a smoother decaying curve than SASRec. Moreover, for SPMRec, the degrees of spectrum smoothness for all four datasets are 26.34 for Beauty, 20.76 for Tools, 20.25 for Toys, and 19.24 for Office. However, SASRec has lower degrees for spectrum smoothness as 19.54 for Beauty, 17.10 for Tools, 17.00 for Toys, and 13.67 for Office. SPMRec consistently has larger degrees of spectrum smoothness than SASRec, which demonstrates the effectiveness of SPMRec in preventing sequences from the degeneration for the sequential recommendation.
6. Related Work
6.1. Sequential Recommendation
Sequential Recommendation (SR) is a technique that encodes the temporal interactions of each user as a sequence and uses the encoded behaviors to infer their preferences. The main idea behind SR is to model the transitions between items within sequences. The earliest SR methods used Markovian approaches, such as FPMC (Rendle et al., 2010) and Fossil (He and McAuley, 2016), with the assumption that the item prediction depends only on the previous interacted items Another line of work utilizes the Recurrent Neural Network (RNN) as the base model for modeling sequential interactions. Several works adopt RNNs to address the SR problem, including GRU4Rec (Hidasi et al., 2015), HGN (Ma et al., 2019; Peng et al., 2021), HRNN (Quadrana et al., 2017), and RRN (Wu et al., 2017). Different variants of RNNs are proposed, such as the original GRU in GRU4Rec and hierarchical RNN in HGN and HRNN. The third line of methods to solve the SR problem is using the Convolution Neural Network (CNN), with the interpretation of a sequence of item embeddings as an image, including Caser (Tang and Wang, 2018), CosRec (Yan et al., 2019), and NextItNet (Yuan et al., 2019). The recent success of the self-attention module (Vaswani et al., 2017) inspires the adaptation of Transformer to the sequential recommendation task. The self-attention module captures the sequential high-order item transitions by modeling attention from all items in the sequence. SASRec (Kang and McAuley, 2018) is the earliest work of applying Transformer to SR. BERT4Rec (Sun et al., 2019) builds upon BERT (Devlin et al., 2019) with bidirectional Transformers for SR. FMLP-Rec (Zhou et al., 2022) proposes an all-MLP architecture in the Transformer for SR. Several variants of Transformers (Fan et al., 2021; Li et al., 2020; Wu et al., 2020; Liu et al., 2021) are further proposed.
6.2. Representation Degeneration
The representation degeneration problem describes the issue of learned embeddings being narrowly distributed in a small cone (Gao et al., 2019). The degeneration issue is firstly proposed in the word representations learning problem. As shown in (Gao et al., 2019), words from text data follow the long-tail distribution, which degenerates the representation learning process. Theoretically, it causes the narrow distribution of words in the representation latent space (Gao et al., 2019). Moreover, an empirical analysis (Yu et al., 2022) shows that the learning of tail words degenerates the representation optimization of frequent words. An important analysis in (Wang et al., 2020) shows that degeneration is closely related to the fast decaying singular values distribution. In other words, the largest singular value is significantly larger than all other singular values.
Distinct from the natural language processing (NLP) area, SR still has several unique characteristics. For example, sequences are short (data sparsity in recommendation systems), while sentences/paragraphs in NLP are typically long. This distinct characteristic introduces additional challenges for SR, where the degeneration issue can also happen in the sequence output embeddings. DuoRec (Qiu et al., 2022) observes that the representation degeneration issue also exists in the self-supervised SR, with data augmentations. DuoRec and DirectAU (Wang et al., 2022b) optimizes the uniformity property in the in-batch representations to enforce the representations to be uniformly distributed, where the uniformity is based on the Euclidean distance. This work connects the item embeddings degeneration with the degeneration of sequence representations. The intertwining of both sequence and item rank degeneration is rarely discussed before in SR. Moreover, some solutions from NLP to the degeneration problem only apply to the item side, including the CosReg (Gao et al., 2019) based on cosine similarities, Euclidean regularization based on learned word similarities (Zhang et al., 2020), spectrum control (Wang et al., 2020; Yan et al., 2022), and gradient gating (Yu et al., 2022).
As discussed in (de Souza Pereira Moreira et al., 2021), natural language processing (NLP) and sequential recommendation (SR) share multiple similarities, including both learning discrete tokens learning (words in NLP and items in SR) and long-tail distribution of frequency for both words and items. With this inspiration, DuoRec (Qiu et al., 2022) observes that the representation degeneration issue also exists in the self-supervised SR, with data augmentations. DuoRec optimizes the uniformity property in the in-batch augmented representations to alleviate the degeneration issue in contrastive learning, where the uniformity is based on the Euclidean distance.